OpenStudy (anonymous):

Find an equation of the tangent line to the graph of the function at the point (1, 1). 1 + ln(xy) = e^(x-y)

OpenStudy (anonymous):

can you help me? :)

OpenStudy (anonymous):

a tangent line to a function is just the derivative of the line, so you need to take the derivative of your function. d/dx(1+ln(xy)-e^(x-y)) (this is slightly rearranged from your initial function) d/dx(1)=0 d/dx(ln(xy))=1/x and d/dx of -e^(x-y)=-e^(x-y) so you have 1/x-e^(x-y) then you need to evaluate it at the point (1,1) so just substitute in your point values into the appropriate variables 1/1-e^(1-1) = 1-e^0 = 1-1 = 0

OpenStudy (anonymous):

hmmm im not sure I am following your work. Also is the equation 0?

OpenStudy (anonymous):

yes, i just switched it over to one side to track things easier myself, sorry for not clarifying. what aren't you following?

OpenStudy (anonymous):

do you know have to take the derivative of something?

OpenStudy (anonymous):

but i do not believe the answer is y=0

OpenStudy (anonymous):

yes i do, but i was confused how to take the derivative of this function

OpenStudy (anonymous):

i could have messed something up, i'm not a math professor let me double check

OpenStudy (anonymous):

haha thank you

OpenStudy (anonymous):

is what you have typed how it's presented to you in the book? is there any other information?

OpenStudy (anonymous):

it is an online homework thing called webassign and that is how the problem was presented

OpenStudy (anonymous):

hmm we use webassign as well unfortunately. i can't really find much in my old notes regarding setups like this unfortunately. hopefully someone else can help you out.

OpenStudy (anonymous):

alright then, thank you for trying

OpenStudy (anonymous):

yeah, sorry i couldn't help.